# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VI-A Experimental setups for personalized machine learning
$\vspace{-0.2em}\min_{\boldsymbol{x}\in\Omega}~{}$ $\textstyle\sum_{i=1}^{m}F_{i}^{t}(x_{i},G^{t}(\boldsymbol{x})),~{}G^{t}(\boldsymbol{x})=\frac{1}{m}\sum_{i=1}^{m}G_{i}^{t}(x_{i}),$
@@
$\textstyle\quad\quad\quad~{}~{}G_{i}^{t}(x_{i})=\frac{1}{|\mathcal{D}_{i}^{t}|}\sum_{j=1}^{|\mathcal{D}_{i}^{t}|}\mathcal{L}(x_{i};\xi_{j}).\vspace{-0.2em}$ (21)
@@
In each experiment, we conducted two types of comparisons: 1) we compared our algorithm with the state-of-the-art distributed aggregative optimization algorithms in [4] and [21] under the same cumulative privacy budget; and 2) to compare with existing DP approaches for distributed optimization, we also implemented the DP approach PDOP in [36] (which uses geometrically decreasing stepsizes and DP-noise variances to ensure a finite cumulative privacy budget) and the DP approach in [48] (which uses time-varying sample sizes to ensure a finite cumulative privacy budget) within our algorithmic framework. For both “MNIST” and “CIFAR-10” datasets, we considered heterogeneous data distributions, where $60\%$ of the data from the $i$th and the $(2i)$th classes was assigned to agent $i$, while the remaining $40\%$ was evenly distributed among the other agents. The projection set in Algorithm 1 and the comparison algorithms was set to $\Omega_{i}=\{x\in\mathbb{R}^{n}|10^{-6}\!\leq\!x_{j}\!\leq\!10^{6},\forall j\!=\!1,\cdots,n\}$ (e.g., with $n=784$ in the “MNIST” experiment). The interaction pattern contains $5$ agents connected in a circle, where each agent can only communicate with its two immediate neighbors. For the weight matrix $W$, we set $w_{ij}\!=\!0.3$ if agents $i$ and $j$ are neighbors, and $w_{ij}\!=\!0$ otherwise.
@@ VI-B Evaluation results on the “MNIST” dataset
In this experiment, we tested the effectiveness of Algorithm 1 using the “MNIST” dataset [64], which contains $60,000$ training images and $10,000$ testing images of handwritten digits. Each image is a $28\times 28$ grayscale image representing a single digit from 00 to $9$. All agents were equipped with the same two-layer convolutional neural network (CNN), consisting of two convolutional layers with $16$ and $32$ filters, with each followed by a max pooling layer. Finally, a fully connected dense layer maps the extracted features to $10$ output classes. We set the stepsizes and DP-noise variances in our algorithm as $\lambda_{x}^{t}=(t+1)^{-0.55}$, $\lambda_{y}^{t}=(t+1)^{-0.02}$, $\lambda_{z}^{t}=(t+1)^{-0.03}$, $\sigma_{i,x}^{t}=0.01(t+1)^{-0.505+0.001i}$, $\sigma_{i,y}^{t}=0.01(t+1)^{-0.005+0.001i}$, and $\sigma_{i,z}^{t}=0.01(t+1)^{-0.015+0.001i}$ for $i=1,\cdots,5$, respectively, which satisfy all conditions in Theorem 1. In our comparison, for the algorithm in [4], we set its stepsize as $\alpha=1$ and its parameter as $\delta=0.5$, and for the algorithm in [21], we set its stepsize as $\lambda_{x}^{t}=(t+1)^{-0.5}$, in line with the guidelines provided in [4] and [21], respectively (see Section IV in [4] and Eq. (8) in [21] for details). Furthermore, the stepsizes and DP-noise variances for the DP approach PDOP in [36] were set to $\lambda_{x}^{t}=0.1(0.94)^{t}$, $\lambda_{y}^{t}=0.1(0.98)^{t}$, $\lambda_{z}^{t}=0.1(0.97)^{t}$, $\sigma_{x}^{t}=0.07(0.99)^{t}$, $\sigma_{y}^{t}=0.008(0.99)^{t}$ and $\sigma_{z}^{t}=0.008(0.99)^{t}$, respectively. For the DP mechanism in [48], the stepsizes, DP-noise variances, decaying factor $\beta^{t}$, and the time-varying sampling number $\gamma_{t}$ were set to $\lambda_{x}^{t}=\frac{0.05}{(t+1)^{0.76}}$, $\lambda_{y}^{t}=\frac{0.05}{(t+1)^{0.11}}$, $\lambda_{z}^{t}=\frac{0.05}{(t+1)^{0.65}}$, $\sigma_{x}^{t}=0.1(t+1)^{0.01}$, $\sigma_{y}^{t}=0.01(t+1)^{0.01}$, $\sigma_{z}^{t}=0.01(t+1)^{0.01}$, $\beta^{t}=\frac{0.05}{(t+1)^{0.51}}$, and $\gamma^{t}=\lceil(t+1)^{1.4}\rceil$, respectively. The parameter settings for [36] and [48] lead to a larger cumulative privacy budget (and hence, weaker privacy protection) than our algorithm, making the optimization-accuracy comparison favorable to these baselines (see the yellow and blue blocks in Fig. 1(c)).
@@
The resulting training loss, test accuracy, and cumulative privacy budget after $5$ epochs of training are summarized in Fig. 1(a), Fig. 1(b), and Fig. 1(c), respectively. It is clear that the proposed algorithm has much lower training loss and higher test accuracy than existing counterparts under LDP constraints. In particular, Fig. 1(a) and Fig. 1(b) show that Laplace noise indeed accumulates in the conventional gradient-tracking-based distributed aggregative optimization algorithms in [4] and [21], leading to divergent behaviors (see the green and gray curves in Fig. 1(a) and Fig. 1(b)). Furthermore, the yellow curves in Fig. 1(a) and Fig. 1(b) indicate that PDOP, which uses geometrically decaying stepsizes and DP-noise variances to ensure a finite cumulative privacy budget in the infinite time horizon, is unable to train the complex CNN model. These comparisons corroborate the advantage of the proposed Algorithm 1.
(a) Training loss
(b) Test accuracy
(c) Cumulative privacy budget of $31,250$ iterations
Figure 2: Comparison of Algorithm 1 with existing distributed aggregative optimization algorithms in [4] and [21] (under the same cumulative privacy budget), and existing DP approaches in [36] and [48] for distributed optimization, using the “CIFAR-10” dataset. The error bars represent the standard deviation.
@@ VI-C Evaluation results on the “CIFAR-10” dataset
In the second experiment, we evaluated the performance of Algorithm 1 by training a four-layer CNN on the “CIFAR-10” dataset [65], which offers greater diversity and complexity compared to the “MNIST” dataset. The dataset contains $50,000$ training images and $10,000$ testing images, with each image being a $32\!\times\!32$ color image belonging to one of $10$ different classes. The CNN consists of four convolutional layers with $32$, $64$, $128$, and $128$ filters, each followed by LeakyReLU activation. Max pooling is applied after the second and fourth layers to reduce spatial dimensions. After global average pooling, the features are fed into a fully connected layer for $10$-class prediction. In this experiment, the stepsizes in our algorithm were set to be the same as those used in the previous “MNIST” experiment, and the DP-noise variances were set to $\sigma_{i,x}^{t}=0.001(t+1)^{-0.505+0.001i}$, $\sigma_{i,y}^{t}=0.001(t+1)^{-0.005+0.001i}$, and $\sigma_{i,z}^{t}=0.001(t+1)^{-0.015+0.001i}$ for all agents, which satisfy all the conditions in Theorem 1. For the comparison algorithms, the noise variances were reduced by a factor of $10$ compared to those used in the “MNIST” experiment, while all other parameters remained unchanged.
The results are summarized in Fig. 2. It can be seen that the proposed algorithm has much better robustness to DP noise compared to existing distributed aggregative optimization algorithms in [4] and [21]. Moreover, compared with the existing DP approaches in [36] and [48] for distributed optimization, our Algorithm 1 achieves lower training loss and higher test accuracy even under a stronger privacy guarantee.
@@ VII Conclusions
To simplify notations, we use $\bar{x}_{(i)}^{t}=\frac{1}{m}(\sum_{l\neq i}x_{l(i)}^{t}+x_{i}^{t})$ to denote the average of agent $i$’s optimization variable $x_{i}^{t}$ and other agents’ estimates $x_{l(i)}$ of this variable for all $l\neq i$. We further define the assembly of the $i$th optimization variable as $\boldsymbol{x}_{(i)}^{t}\!\!=\!\!\mathrm{col}(x_{1(i)}^{t},\cdots,x_{i}^{t},\cdots,x_{m(i)}^{t})\!\in\!\mathbb{R}^{mn_{i}}$. According to the definition of $R_{i}\!\!=\!\![\boldsymbol{0}_{n_{i}\times n_{1}},\cdots,I_{n_{i}},\cdots,\boldsymbol{0}_{n_{i}\times n_{m}}]$, we have $R_{i}\boldsymbol{x}_{i}^{t}=x_{i}^{t}$, and hence, we do not distinguish between them in this section. To further simplify the notations, we define $f(\boldsymbol{x},\boldsymbol{y})=\sum_{i=1}^{m}f_{i}(x_{i},y_{i})\in\mathbb{R}$, $g(\boldsymbol{x})=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x_{i})\in\mathbb{R}^{r}$, $\nabla_{x}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})\!=\!\mathrm{col}(\nabla_{x}f_{1}(x_{1},y_{1}),\cdots,\nabla_{x}f_{m}(x_{m},y_{m}))\in\mathbb{R}^{n}$, $\nabla_{y}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})\!=\!\mathrm{col}(\nabla_{y}f_{1}(x_{1},y_{1}),\cdots,\nabla_{y}f_{m}(x_{m},y_{m}))\in\mathbb{R}^{mr}$, $\boldsymbol{g}(\boldsymbol{x})\!=\!\mathrm{col}(g_{1}(x_{1}),\cdots,g_{m}(x_{m}))\in\mathbb{R}^{mr}$, and $\nabla\boldsymbol{g}(\boldsymbol{x})=\text{blkdiag}(\nabla g_{1}(x_{1}),\cdots,\nabla g_{m}(x_{m}))\in\mathbb{R}^{n\times mr}$. Moreover, we introduce some additional notations: $u_{i}^{t}\!=\!\nabla_{x}f_{i}^{t}(x_{i}^{t},\tilde{y}_{i}^{t})-\nabla g_{i}^{t}(x_{i}^{t})\tilde{z}_{i}^{t}$, $\hat{x}_{(i)}^{t}=x_{i}^{t}-\bar{x}_{(i)}^{t}$, $\hat{y}_{i}^{t}=y_{i}^{t}-\bar{y}^{t}$, $\hat{z}_{i}^{t}=z_{i}^{t}-\bar{z}^{t}$, $L_{f}=\max\{L_{f,1},L_{f,2}\}$, and $\bar{L}_{f}=\max\{\bar{L}_{f,1},\bar{L}_{f,2}\}$.
@@ VII-A Auxiliary lemmas
In this section, we introduce some auxiliary lemmas that will be used in our subsequent convergence analysis.
@@
Under Assumption 1, $\nabla F(\boldsymbol{x})$ is $L_{F}$-Lipschitz continuous, i.e., for any $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in\Omega$, we have
$\|\nabla F(\boldsymbol{x}_{1})-\nabla F(\boldsymbol{x}_{2})\|\leq L_{F}\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2}\|,$
with $L_{F}=\bar{L}_{f,1}+L_{f}\bar{L}_{g}+L_{g}(\bar{L}_{f,1}+\bar{L}_{f,2}+L_{g}\bar{L}_{f,2})$.
@@
For real numbers $a,b,c,d>0$ satisfying $\frac{c}{d}<\frac{d}{b}$, the inequality $\frac{d}{b}<\frac{c+d}{a+b}<\frac{c}{a}$ always holds. Hence, for any $t\in\mathbb{N}^{+}$ and $p\in[0,t)$, we have the inequality $\frac{1}{(t-p)^{2}}<(\frac{p+1}{t})^{2}$ by setting $a=p$, $b=t-p$, $c=p$, and $d=1$.
By using Lemma 3 and the relationship $\frac{1}{(t-p)^{2}}<(\frac{p+1}{t})^{2}<(\frac{p+1}{t})^{v}$ for any $p\in[0,t)$ and $v\in(0,2)$, we obtain
@@
with the constant $a=4^{-1}(\ln(\sqrt{\gamma})e)^{2}$.
Using (22) and the geometric series formula, we arrive at
$\sum_{p=0}^{t}\frac{\gamma^{t-p}}{(p+1)^{v}}\leq\sum_{p=0}^{t}\sqrt{\gamma}^{t-p}\frac{1}{at^{v}}\leq\frac{4\times 2^{v}}{(\ln(\sqrt{\gamma})e)^{2}(1-\sqrt{\gamma})(t+1)^{v}}.$
@@
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t+1}\|^{2}]$ (23)
$\textstyle\leq(1-|\rho_{2}|)\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$
$\textstyle\quad+\frac{c_{x1}(\lambda_{x}^{t})^{2}}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\boldsymbol{\hat{z}}^{t+1}-\boldsymbol{\hat{z}}^{t}\|^{2}]+c_{x2}(\sigma_{x}^{t})^{2},$

### v3

@@ VI-A Experimental setups for personalized machine learning
$\vskip-1.99997pt\min_{\boldsymbol{x}\in\Omega}$~ $\textstyle\sum_{i=1}^{m}F_{i}^{t}(x_{i},G^{t}(\boldsymbol{x})),~G^{t}(\boldsymbol{x})=\frac{1}{m}\sum_{i=1}^{m}G_{i}^{t}(x_{i}),$
@@
$\textstyle\quad\quad\quad~~G_{i}^{t}(x_{i})=\frac{1}{|\mathcal{D}_{i}^{t}|}\sum_{j=1}^{|\mathcal{D}_{i}^{t}|}\mathcal{L}(x_{i};\xi_{j}).\vskip-1.99997pt$ (22)
@@
In each experiment, we conducted two types of comparisons: 1) we compared our algorithm with the state-of-the-art distributed aggregative optimization algorithms in [6] and [36] under the same cumulative privacy budget; and 2) to compare with existing DP approaches for distributed optimization, we also implemented the DP approach PDOP in [26] (which uses geometrically decreasing stepsizes and DP-noise variances to ensure a finite cumulative privacy budget) and the DP approach in [51] (which uses time-varying sample sizes to ensure a finite cumulative privacy budget) within our algorithmic framework. Note that the first type of comparison evaluates the performance of our novel distributed aggregative optimization algorithm structure, whereas the second type of comparison assesses the performance of our proposed LDP mechanism. For both “MNIST” and “CIFAR-10” datasets, we considered heterogeneous data distributions, where $60\%$ of the data from the $i$th and the $(2i)$th classes was assigned to agent $i$, while the remaining $40\%$ was evenly distributed among the other agents. The parameter $\lambda$ in (22) was set to $\lambda\!=\!0.5$ for the “MNIST” experiment and to $\lambda\!=\!0.1$ for the “CIFAR-10” experiment. The projection set in Algorithm 1 and the comparison algorithms was set to $\Omega_{i}=\{x\in\mathbb{R}^{n}|10^{-6}\!\leq\!x_{j}\!\leq\!10^{6},\forall j\!=\!1,\cdots,n\}$ (e.g., with $n=784$ in the “MNIST” experiment). The interaction pattern contains $5$ agents connected in a circle, where each agent can only communicate with its two immediate neighbors. For the weight matrix $W$, we set $w_{ij}\!=\!0.3$ if agents $i$ and $j$ are neighbors, and $w_{ij}\!=\!0$ otherwise.
(a) Comparison of different algorithms
(b) Cumulative privacy budgets
(c) Comparison under different $\delta$
(d) Comparison under different $\lambda$
Figure 2: Experimental results on CNN training using the “CIFAR-10” dataset. (a) and (b) Training loss and cumulative privacy budget comparison of Algorithm 1 with existing distributed aggregative optimization algorithms in [6] and [36] under the same cumulative privacy budget, and existing DP approaches in [26] and [51] for distributed optimization. (c) Training losses of Algorithm 1 under different $\delta$ ($\delta$ is given in Corollary 1). (d) Training losses of Algorithm 1 under different $\lambda$ ($\lambda$ is defined in (22)). The error bars represent the standard deviation.
@@ VI-B Evaluation results on the “MNIST” dataset
In this experiment, we tested the effectiveness of Algorithm 1 using the “MNIST” dataset [30], which contains $60,000$ training images and $10,000$ testing images of handwritten digits. Each image is a $28\times 28$ grayscale image representing a single digit from 0 to $9$. All agents were equipped with the same two-layer convolutional neural network (CNN), consisting of two convolutional layers with $16$ and $32$ filters, with each followed by a max pooling layer. Finally, a fully connected dense layer maps the extracted features to $10$ output classes. We set the stepsizes and DP-noise variances in our algorithm as $\lambda_{x}^{t}=(t+1)^{-0.55}$, $\lambda_{y}^{t}=(t+1)^{-0.02}$, $\lambda_{z}^{t}=(t+1)^{-0.03}$, $\sigma_{i,x}^{t}=0.01(t+1)^{-0.505+0.001i}$, $\sigma_{i,y}^{t}=0.01(t+1)^{-0.005+0.001i}$, and $\sigma_{i,z}^{t}=0.01(t+1)^{-0.015+0.001i}$ for $i=1,\cdots,5$, respectively, which satisfy all conditions in Theorem 1. In our comparison, for the algorithm in [6], we set its stepsize as $\alpha=1$ and its parameter as $\delta=0.5$, and for the algorithm in [36], we set its stepsize as $\lambda_{x}^{t}=(t+1)^{-0.5}$, in line with the guidelines provided in [6] and [36], respectively (see Section IV in [6] and Eq. (8) in [36] for details). Furthermore, the stepsizes and DP-noise variances for the DP approach PDOP in [26] were set to $\lambda_{x}^{t}=0.1(0.94)^{t}$, $\lambda_{y}^{t}=0.1(0.98)^{t}$, $\lambda_{z}^{t}=0.1(0.97)^{t}$, $\sigma_{x}^{t}=0.07(0.99)^{t}$, $\sigma_{y}^{t}=0.008(0.99)^{t}$ and $\sigma_{z}^{t}=0.008(0.99)^{t}$, respectively. For the DP mechanism in [51], the stepsizes, DP-noise variances, decaying factor $\beta^{t}$, and the time-varying sampling number $\gamma_{t}$ were set to $\lambda_{x}^{t}=\frac{0.05}{(t+1)^{0.76}}$, $\lambda_{y}^{t}=\frac{0.05}{(t+1)^{0.11}}$, $\lambda_{z}^{t}=\frac{0.05}{(t+1)^{0.65}}$, $\sigma_{x}^{t}=0.1(t+1)^{0.01}$, $\sigma_{y}^{t}=0.01(t+1)^{0.01}$, $\sigma_{z}^{t}=0.01(t+1)^{0.01}$, $\beta^{t}=\frac{0.05}{(t+1)^{0.51}}$, and $\gamma^{t}=\lceil(t+1)^{1.4}\rceil$, respectively. The parameter settings for [26] and [51] lead to a larger cumulative privacy budget (and hence, weaker privacy protection) than our algorithm, making the optimization-accuracy comparison favorable to these baselines (see the yellow and blue blocks in Fig. 1(b)).
@@
The experimental results are summarized in Fig. 1. It is clear that the proposed algorithm has much lower training loss than existing counterparts under LDP constraints. In particular, Fig. 1(a) shows that Laplace noise indeed accumulates in the conventional gradient-tracking-based distributed aggregative optimization algorithms in [6] and [36], leading to divergent behaviors (see the green and gray curves in Fig. 1(a)). Furthermore, the yellow curve in Fig. 1(a) indicate that PDOP, which uses geometrically decaying stepsizes and DP-noise variances to ensure a finite cumulative privacy budget in the infinite time horizon, is unable to train the complex CNN model. These comparisons corroborate the advantage of the proposed Algorithm 1. Fig. 1(c) shows that our algorithm is robust even when $\delta$ slightly deviates from the range $\delta\in(0,\frac{1}{14})$ derived in Corollary 1. Moreover, a smaller $\delta$ leads to lower training loss. Fig. 1(d) shows that an overly small or large $\lambda$ results in high training loss, whereas a moderate $\lambda$ improves performance.
@@ VI-C Evaluation results on the “CIFAR-10” dataset
In the second experiment, we evaluated the performance of Algorithm 1 by training a four-layer CNN on the “CIFAR-10” dataset [29], which offers greater diversity and complexity compared to the “MNIST” dataset. The dataset contains $50,000$ training images and $10,000$ testing images, with each image being a $32\!\times\!32$ color image belonging to one of $10$ different classes. The CNN consists of four convolutional layers. Max pooling is applied after the second and fourth layers to reduce spatial dimensions. After global average pooling, the features are fed into a fully connected layer for $10$-class prediction. In this experiment, the stepsizes in our algorithm were set to be the same as those used in the previous “MNIST” experiment, and the DP-noise variances were set to $\sigma_{i,x}^{t}=0.001(t+1)^{-0.505+0.001i}$, $\sigma_{i,y}^{t}=0.001(t+1)^{-0.005+0.001i}$, and $\sigma_{i,z}^{t}=0.001(t+1)^{-0.015+0.001i}$ for all agents, which satisfy all the conditions in Theorem 1. For the comparison algorithms, the noise variances were reduced by a factor of $10$ compared to those used in the “MNIST” experiment, while all other parameters remained unchanged.
The results are summarized in Fig. 2. It can be seen that the proposed algorithm has much better robustness to DP noise compared to existing distributed aggregative optimization algorithms in [6] and [36]. Moreover, compared with the existing DP approaches in [26] and [51] for distributed optimization, our Algorithm 1 achieves lower training loss even under a stronger privacy guarantee. The trends in the sensitivity analysis shown in Fig. 2(c) and Fig. 2(d) closely resemble those in Fig. 1(c) and Fig. 1(d), respectively.
@@ VII Conclusions
To simplify notations, we use $\bar{\mathrm{x}}_{(i)}^{t}=\frac{1}{m}(\sum_{l\neq i}\mathrm{x}_{l(i)}^{t}+x_{i}^{t})$ to denote the average of agent $i$’s optimization variable $x_{i}^{t}$ and other agents’ estimates $\mathrm{x}_{l(i)}^{t}$ of this variable for all $l\neq i$. We further define the assembly of the $i$th optimization variable as $\mathbf{x}_{(i)}^{t}\!\!=\!\!\mathrm{col}(\mathrm{x}_{1(i)}^{t},\cdots,\mathrm{x}_{m(i)}^{t})\!\in\!\mathbb{R}^{mn_{i}}$. To further simplify the notations, we define $f(\boldsymbol{x},\boldsymbol{y})=\sum_{i=1}^{m}f_{i}(x_{i},y_{i})\in\mathbb{R}$, $g(\boldsymbol{x})=\frac{1}{m}\sum_{i=1}^{m}g_{i}(x_{i})\in\mathbb{R}^{r}$, $\nabla_{x}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})\!=\!\mathrm{col}(\nabla_{x}f_{1}(x_{1},y_{1}),\cdots,\nabla_{x}f_{m}(x_{m},y_{m}))\in\mathbb{R}^{n}$, $\nabla_{y}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y})\!=\!\mathrm{col}(\nabla_{y}f_{1}(x_{1},y_{1}),\cdots,\nabla_{y}f_{m}(x_{m},y_{m}))\in\mathbb{R}^{mr}$, $\boldsymbol{g}(\boldsymbol{x})\!=\!\mathrm{col}(g_{1}(x_{1}),\cdots,g_{m}(x_{m}))\in\mathbb{R}^{mr}$, and $\nabla\boldsymbol{g}(\boldsymbol{x})=\text{blkdiag}(\nabla g_{1}(x_{1}),\cdots,\nabla g_{m}(x_{m}))\in\mathbb{R}^{n\times mr}$. Moreover, we introduce some additional notations: $u_{i}^{t}\!=\!\nabla_{x}f_{i}^{t}(x_{i}^{t},\tilde{y}_{i}^{t})+\nabla g_{i}^{t}(x_{i}^{t})\tilde{z}_{i}^{t}$, $\hat{x}_{(i)}^{t}=x_{i}^{t}-\bar{\mathrm{x}}_{(i)}^{t}$, $\hat{y}_{i}^{t}=y_{i}^{t}-\bar{y}^{t}$, $\hat{z}_{i}^{t}=z_{i}^{t}-\bar{z}^{t}$, $L_{f}=\max\{L_{f,1},L_{f,2}\}$, and $\bar{L}_{f}=\max\{\bar{L}_{f,1},\bar{L}_{f,2}\}$.
@@ VII-A Auxiliary lemmas
Under Assumption 1, $\nabla F(\boldsymbol{x})$ is $L_{F}$-Lipschitz continuous, i.e., for any $\boldsymbol{x}_{1},\boldsymbol{x}_{2}\in\Omega$, we have $\|\nabla F(\boldsymbol{x}_{1})-\nabla F(\boldsymbol{x}_{2})\|\leq L_{F}\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2}\|,$ with $L_{F}=\bar{L}_{f,1}+L_{f}\bar{L}_{g}+L_{g}(\bar{L}_{f,1}+\bar{L}_{f,2}+L_{g}\bar{L}_{f,2})$.
@@
For real numbers $a,b,c,d>0$ satisfying $\frac{d}{b}\leq\frac{c}{a}$, the inequality $\frac{d}{b}\leq\frac{c+d}{a+b}\leq\frac{c}{a}$ always holds. Hence, for any $t\in\mathbb{N}^{+}$ and $p\in(0,t)$, we have the inequality $\frac{1}{(t-p)^{2}}\leq(\frac{p+1}{t})^{2}$ by setting $a=c=p$, b=tpb=t-p, and $d=1$.
By using Lemma 3 and the relationship $\frac{1}{(t-p)^{2}}\leq(\frac{p+1}{t})^{2}<(\frac{p+1}{t})^{v}$ for any $p\in(0,t)$ and $v\in(0,2)$, we obtain
@@
with $a\!=\!4^{-1}(\ln(\sqrt{\gamma})e)^{2}$. Note that (23) still holds when p=0p\!=\!0.
Using (23) and the geometric series formula, we arrive at $\sum_{p=0}^{t}\frac{\gamma^{t-p}}{(p+1)^{v}}\leq\sum_{p=0}^{t}\sqrt{\gamma}^{t-p}\frac{1}{at^{v}}\leq\frac{4\times 2^{v}}{(\ln(\sqrt{\gamma})e)^{2}(1-\sqrt{\gamma})(t+1)^{v}}.$
@@
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}]\leq\frac{c_{x1}(\lambda_{x}^{t})^{2}}{(\lambda_{z}^{t})^{2}}\mathbb{E}[\|\boldsymbol{\hat{z}}^{t+1}-\boldsymbol{\hat{z}}^{t}\|^{2}]$
$\textstyle+(1-|\rho_{2}|)\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]\!+\!c_{x2}(\sigma_{x}^{t})^{2},\vskip-1.99997pt$ (24)